OpenStudy (anonymous):

integral of 1/( x^2 +6x +9) dx

OpenStudy (anonymous):

does it help to know that \(x^2+6x+9=(x+3)^2\)?

OpenStudy (experimentx):

(x+3)^-2

OpenStudy (ash2326):

\[\int \frac{1}{x^2+6x+9} dx\] If you notice \[x^2+6x+9\] this is \[(x+3)^2\] so we have \[\int \frac{1}{(x+3)^2} dx\] or \[\int (x+3)^{-2} dx\] Can you solve now?

OpenStudy (anonymous):

i guess it does!

OpenStudy (anonymous):

YEs thanks

OpenStudy (anonymous):

its -1/(x+3)+c

OpenStudy (ash2326):

Yeah great work:D